solve the following system of equations. y=(y-2)^2 -4 y= -1/2
Answer by stanbon(75887)

solve the following system of equations.
y = (y-2)^2 -4
y= -1/2
========
Supposedly, y = -1/2
---
Checking in the 1st equation::
-1/2 = (-1/2 - 2)^2 -4
-------
-1/2 = (-5/2)^2 - 4
-1/2 = (25/4) - (16/4)
-1/2 = 9/4
That is a contradiction.
----
Please check your post.
Cheers,
Stan H.